# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 3.2 Formulation of model-oriented distance via posets
Definition 2 provides a succinct yet principled strategy for defining a model-oriented distance across a wide range of graph classes. However, since the definition is implicit, it is not immediately clear how to compute the distance for any specific class. We now provide a constructive characterization of the model-oriented distance which we both use to study under what conditions it simplifies (see Sections 3.3 and 4) and to develop algorithmic tools for computing it (see Section 5).
@@
A poset (partially ordered set) $\mathcal{L}:=(L,\preceq)$ is a collection $L$ of elements and a relation $\preceq$ that is reflexive ($x\preceq x,~\forall x\in L$), transitive ($x\preceq y,y\preceq z\Rightarrow x\preceq z,~\forall x,y,z\in L$), and anti-symmetric ($x\preceq y,y\preceq x\Rightarrow x=y,~\forall x,y\in L$). We use notation $x\prec y$ if $x\preceq y$ and $x\neq y$. An element $y\in L$ covers $x\in L$, denoted as $x\lessdot y$, if $x\prec y$ and there is no $z\in L$ with $x\prec z\prec y$. The poset $\mathcal{L}$ is connected if for every $x,y\in L$, there exists a finite sequence of $k\in\mathbb{N}_{+}$ elements $x=x_{1},x_{2},\dots,x_{k}=y$ such that either $x_{i+1}\gtrdot x_{i}$ or $x_{i+1}\lessdot x_{i}$ for every $i=1,2,\dots,k-1$. An element $x\in L$ is called the least (greatest) element if $x\preceq y$ ($x\succeq y$) for all $y\in L$. Elements $x,y\in\mathcal{L}$ are said to be comparable if either $x\preceq y$ or $y\succeq x$; otherwise, they are incomparable.
@@
Given a graph class $\mathfrak{G}$ and an associated model map $\mathcal{M}$ that satisfy Condition˜1, the corresponding model-oriented poset is $\mathcal{L}:=(\mathfrak{G},\preceq)$ with $\preceq$’ given by model containment: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we say $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
@@
Reflexivity and transitivity of $\preceq$’ hold by definition, while the anti-symmetry follows from the injectivity of $\mathcal{M}(\cdot)$ (Condition˜1). The poset $\mathcal{L}$ offers a natural way to organize the graphs $\mathcal{G}\in\mathfrak{G}$ based on the models they represent. The neighbors in Definition˜1 are exactly the covering relations in the poset: $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if either $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ or $\mathcal{G}_{1}$ covers $\mathcal{G}_{2}$. Note that Condition˜2 is equivalent to requiring that the poset $\mathcal{L}$ is connected, which is typically ensured by the existence of a least element $\mathcal{G}_{\hat{0}}$ (usually the empty graph) in $\mathfrak{G}$.
@@
Figure 3: Hasse diagrams for different statistical graphs over vertex set $\{1,2,3\}$: each box represents a graph, and each dashed arrow represents a covering relation, i.e., we draw $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$ pointing upwards if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$. We can determine whether $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ by checking whether there is a directed path from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$. In (b) and (d), only part of the poset is shown for simplicity. Each poset has a least element, but not all have a greatest element. The poset in panel (d) is not graded when the graphs have more than three vertices (see Proposition˜10). This means there is no consistent way to assign ranks to the graphs.
@@
For example, consider $\mathfrak{G}$ to be all the UGs over the vertex set $\{1,2,3\}$ and let $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ be the model map. In this case, the partial order $\preceq$’ can be characterized graphically: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, one can show that $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$.
@@
Further, $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ if and only if $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ and $\mathcal{G}_{2}$ has exactly one more edge than $\mathcal{G}_{1}$. The least element $\mathcal{G}_{\hat{0}}$ in the poset is the empty graph. This poset can be visualized with a Hasse diagram shown in Fig.˜3(a): each box is a a graph; if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$, we then place $\mathcal{G}_{2}$ above $\mathcal{G}_{1}$ and draw an upward arrow $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$. Partial order can be read off from the diagram: $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if there exists a directed path (can be of length zero) from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$. Hence, the poset $\mathcal{L}=(\mathfrak{G},\preceq)$ can be compactly represented by its Hasse diagram, which is directed acyclic graph over $\mathfrak{G}$. Finally, it is not hard to see that our model-oriented distance, denoted hereafter as $d_{\mathcal{L}}$ to highlight its dependence on $\mathcal{L}$ and hence $(\mathfrak{G},\mathcal{M})$, is given by
@@
where the path can contain dashed arrows in both directions.
We note that the ‘shortest path length’ is a standard way to measure distance in posets (Monjardet, 1981; Trotter, 1992; Brualdi, Graves and Lawrence, 1995), although its use has primarily been confined to the order theory literature.
3.3 Poset structures that facilitate the computation of model-oriented distance
Having established that the model-oriented distance corresponds to the shortest-path distance in the Hasse diagram of the model-oriented poset, we can compute it by using using the well-known $A^{*}$ algorithm. In general, however, this can be expensive, particularly because it requires repeatedly identifying all neighbors of a graph while traversing many candidate paths in the Hasse diagram. Often, however, we can exploit additional structure in the poset to simplify computation — either by deriving a closed-form expression for the distance (eliminating the need for $A^{*}$), by streamlining the neighborhood search, or by reducing the number of paths that must be considered. In the following, we study several such structural properties in general, as we will rely on them when analyzing the model-oriented distances corresponding to probabilistic UG and CPDAG models, as well as causal DAGs and MPDAG models (see Section 4). We also examine the conditions under which the model-oriented distance reduces to an edge-by-edge distance similarly to the SHD.
@@ 3.3.1 Graded poset
A connected poset $\mathcal{L}$ is graded if there exists a function $\operatorname{\mathrm{rank}}:\mathfrak{G}\rightarrow\mathbb{N}$ such that $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ implies $\operatorname{\mathrm{rank}}(\mathcal{G}_{2})=\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1$ and $\operatorname{\mathrm{rank}}(\mathcal{G})=0$ for a minimal $\mathcal{G}\in\mathfrak{G}$. It can be shown that if it exists, the rank function is unique. A graph $\mathcal{G}^{\prime}\in\mathfrak{G}$ is called a minimal element if there exists no $\mathcal{G}\in\mathfrak{G}$ such that $\mathcal{G}\prec\mathcal{G}^{\prime}$ ($\mathcal{G}\succ\mathcal{G}^{\prime}$).
@@
Proposition 1.
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is a graded poset with rank function $\mathrm{rank}(\cdot)$. For any $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ with $\mathcal{G}_{s}\succeq\mathcal{G}_{t}$, every downward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ has length $\mathrm{rank}(\mathcal{G}_{s})-\mathrm{rank}(\mathcal{G}_{t})$, which also equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
This is a well-known result that follows from the property that in a graded poset every maximal chain that is bounded between two comparable elements are of the same length. A subset of $\mathfrak{G}$ is a chain (i.e., totally ordered set) if every pair of elements are comparable. A chain is maximal if it is not contained by a larger chain. For completeness, we provide a short proof in Section˜C.1.
Second, gradedness makes it significantly easier to enumerate all neighbors of a given graph $\mathcal{G}$. Recall that $\mathcal{G}^{\prime}$ is a neighbor of $\mathcal{G}$ in $\mathcal{L}$ if either $\mathcal{G}^{\prime}\preceq\mathcal{G}$ and there is no distinct graph $\tilde{\mathcal{G}}$ with $\mathcal{G}^{\prime}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}$, or $\mathcal{G}\preceq\mathcal{G}^{\prime}$ and there is no distinct graph $\tilde{\mathcal{G}}$ with $\mathcal{G}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}^{\prime}$. In general, enumerating all neighbors can be cumbersome, since one must check for the absence of these intermediate elements. If $\mathcal{L}$ is graded, this process becomes much simpler since two graphs can only be neighbors if their ranks differ by exactly one. Combined with the requirement that one graph is either above or below the other in the poset, this condition becomes both necessary and sufficient. This characterization is particularly useful when the rank function has a simple form. For instance, we will show in Section 4.1 that the poset underlying undirected graphs (UGs) is graded with the rank of a graph equal to its number of edges. All neighbors of a UG $\mathcal{G}$ are therefore UGs $\mathcal{G}^{\prime}$ that differ by exactly one edge and are comparable to $\mathcal{G}$. In a non-graded poset, neighboring graphs may differ more substantially; for example, we show in Section 4.4 that for causal MPDAGs, a neighboring graph may have many more edges.
@@ 3.3.2 Down-up and up-down distance, semi-lattice, semimodularity
To reduce the number of paths we need to explore, it is natural to consider the shortest path among a restricted set of paths between two graphs in the Hasse diagram. Specifically, for $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we consider the set of down-up and up-down paths between them, denoted by $\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ and $\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ respectively. A path is down-up if it takes the form of
@@
where $i^{\ast}\in\{0,\dots,d\}$ is the inflection point; chains ($i^{\ast}=0$ or $i^{\ast}=d$) are special cases of down-up paths. Conversely, an up-down path takes the form of
@@
Proposition 2.
The set of down-up paths $\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a least element (Condition˜3). Similarly, the set of up-down paths $\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a greatest element (Condition˜3).
@@
The posets of all graph types (UGs, DAGs, CPDAGs, and MPDAGs) we consider in this paper have the empty graph as the least element. Further, except for DAGs, all of them also have a greatest element. See Table˜1 for a summary.
The down-up and up-down paths lead to the definition of down-up and up-down distances.
@@
Under Condition˜3, the down-up distance is defined as $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. Under Condition˜3, the up-down distance is defined as $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
@@
Proposition 3.
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\mathrm{rank}(\cdot)$. Then, under Condition˜3, we have
@@
Any down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ moves downward from $\mathcal{G}_{s}$, and then upward from an inflection point $\mathcal{G}$ to $\mathcal{G}_{t}$. Since $\mathcal{G}\preceq\mathcal{G}_{s}$ and $\mathcal{G}\preceq\mathcal{G}_{t}$, Proposition 1 implies that the length of the path is $\mathrm{rank}(\mathcal{G}_{s})+\mathrm{rank}(\mathcal{G}_{t})-2\mathrm{rank}(\mathcal{G})$. Therefore, the shortest down-up path corresponds to an inflection point $\mathcal{G}$ that maximizes $\mathrm{rank}(\mathcal{G})$ over all $\mathcal{G}\preceq\mathcal{G}_{s},\mathcal{G}\preceq\mathcal{G}_{t}$. The proof for the up-down distance follows similarly. ∎
The following additional simplification requires notions of meet and join semi-lattice structures. In particular, graph $\mathcal{G}_{3}\in\mathfrak{G}$ is the least upper bound (or join) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, denoted as $\mathcal{G}_{1}\vee\mathcal{G}_{2}$, if $\mathcal{G}_{3}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$ and $\mathcal{G}_{3}\preceq\mathcal{G}^{\prime}$ for every $\mathcal{G}^{\prime}$ satisfying $\mathcal{G}^{\prime}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$. The greatest lower bound (or meet) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is defined similarly and is denoted as $\mathcal{G}_{1}\wedge\mathcal{G}_{2}$. Poset $\mathcal{L}$ is a join (meet) semi-lattice if $\mathcal{G}\vee\mathcal{G}^{\prime}$ ($\mathcal{G}\wedge\mathcal{G}^{\prime}$) exists for every $\mathcal{G},\mathcal{G}^{\prime}\in\mathfrak{G}$.
Corollary 1.
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\mathrm{rank}(\cdot)$. Under Condition˜3, if $\mathcal{L}$ is also a meet semi-lattice, we have
@@
We prove Theorem˜2 in Appendix˜D. Fig.˜4 illustrates the basic idea: when $\mathcal{L}$ is semimodular, any shortest path can be transformed into a down-up or up-down path with the same length. We remark that, to our best knowledge, our characterization weakens the conditions for existing results in the literature and is well-suited for studying model-oriented graph distances. For example, Monjardet (1981, Theorem 8) additionally requires the poset to be graded, while Foldes and Radeleczki (2021, Proposition 2.5) instead requires the poset to be a semi-lattice. However, the model-oriented posets for graphs may not satisfy these conditions. In fact, in Section˜4.3 we will use Theorem˜2 to study polytree CPDAGs, where these earlier results are inapplicable.

### v2

@@ 3.2 Formulation of model-oriented distance via posets
Definition 2 provides a succinct yet principled strategy for defining a model-oriented distance across a wide range of graph classes. However, since the definition is implicit, it is not immediately clear how to compute the distance for any specific class. We use partially ordered sets to construct the model-oriented distance, which leads to structural properties that facilitate distance computation (see Sections˜3.3 and 4) and efficient algorithms (see Section˜5).
@@
A poset (partially ordered set) $\mathcal{L}:=(\mathrm{L},\preceq)$ is a collection $\mathrm{L}$ of elements and a relation $\preceq$ that is reflexive ($x\preceq x,~\forall x\in\mathrm{L}$), transitive ($x\preceq y,y\preceq z\Rightarrow x\preceq z,~\forall x,y,z\in\mathrm{L}$), and anti-symmetric ($x\preceq y,y\preceq x\Rightarrow x=y,~\forall x,y\in\mathrm{L}$). We use notation $x\prec y$ if $x\preceq y$ and $x\neq y$. An element $y\in\mathrm{L}$ covers $x\in\mathrm{L}$, denoted as $x\lessdot y$, if $x\prec y$ and there is no $z\in\mathrm{L}$ with $x\prec z\prec y$. The poset $\mathcal{L}$ is connected if for every $x,y\in\mathrm{L}$, there exists a sequence of $k$ elements $x=x_{1},x_{2},\dots,x_{k}=y$ such that either $x_{i+1}\gtrdot x_{i}$ or $x_{i+1}\lessdot x_{i}$ for every $i=1,2,\dots,k-1$. An element $x\in\mathrm{L}$ is a minimal (maximal) element if there is no $y\in\mathrm{L}$ such that $y\prec x$ ($y\succ x$). An element $x\in\mathrm{L}$ is called the least (greatest) element if $x\preceq y$ ($x\succeq y$) for all $y\in\mathrm{L}$. Elements $x,y\in\mathcal{L}$ are said to be comparable if either $x\preceq y$ or $y\succeq x$; otherwise, they are incomparable.
@@
Given a graph class $\mathfrak{G}$ and an associated model map $\mathcal{M}$ that satisfies Condition˜1, the corresponding model-oriented poset is $\mathcal{L}:=(\mathfrak{G},\preceq)$ with $\preceq$’ given by model containment: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we say $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
@@
Reflexivity and transitivity of $\preceq$’ hold by definition, while the anti-symmetry follows from the injectivity of $\mathcal{M}(\cdot)$ (Condition˜1). The poset $\mathcal{L}$ offers a natural way to organize the graphs $\mathcal{G}\in\mathfrak{G}$ based on the models they represent. The neighbors in Definition˜1 are exactly given by the covering relations: $\mathcal{G}^{\prime}$ is a neighbor of $\mathcal{G}$ if either $\mathcal{G}$ covers $\mathcal{G}^{\prime}$ or $\mathcal{G}^{\prime}$ covers $\mathcal{G}$, i.e.,
$\text{Neighbors}(\mathcal{G})=\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\lessdot\mathcal{G}\}\cup\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\gtrdot\mathcal{G}\},\quad\mathcal{G}\in\mathfrak{G}.$ (5)
Note that Condition˜2 is equivalent to requiring that the poset $\mathcal{L}$ is connected, which is typically ensured by the existence of a least element $\mathcal{G}_{\hat{0}}$ (usually the empty graph) in $\mathfrak{G}$.
@@
Figure 3: Hasse diagrams for four different statistical graph classes over vertex set $\{1,2,3\}$: each box represents a graph, and each dashed arrow represents a covering relation, i.e., we draw $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$ pointing upwards if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$. We can determine whether $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ by checking whether there is a directed path from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$. In (b) and (d), only part of the poset is shown for simplicity. Each poset has a least element; posets (a), (c) and (d) also have a greatest element. Posets (a), (b) and (c) and graded; (d) is not graded (see Proposition˜8).
@@
For example, consider $\mathfrak{G}$ to be all the UGs over the vertex set $\{1,2,3\}$ and let $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ be the model map. In this case, the partial order $\preceq$’ can be characterized graphically: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, one can show that $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$. Further, $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ if and only if $\mathcal{G}_{2}$ has exactly one more edge than $\mathcal{G}_{1}$. The least element $\mathcal{G}_{\hat{0}}$ in the poset is the empty graph.
@@
This poset can be visualized with a Hasse diagram shown in Fig.˜3(a): each box is a graph; if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$, we then place $\mathcal{G}_{2}$ above $\mathcal{G}_{1}$ and draw an upward arrow $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$. Partial order can be read off from the diagram: $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if there exists a directed path (can be of length zero) from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$. Hence, the poset $\mathcal{L}=(\mathfrak{G},\preceq)$ can be compactly represented by its Hasse diagram, which is directed acyclic graph over $\mathfrak{G}$. Finally, it is not hard to see that our model-oriented distance, denoted hereafter as $d_{\mathcal{L}}$ to highlight its dependence on $\mathcal{L}$ and hence $(\mathfrak{G},\mathcal{M})$, is given by
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\min\left\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\right\},\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G},$
where
@@
$=\left\{(\mathcal{G}_{0},\dots,\mathcal{G}_{d}):\mathcal{G}_{0}=\mathcal{G}_{s},\mathcal{G}_{d}=\mathcal{G}_{t},\,\text{and for $1\leq i\leq d$, $\mathcal{G}_{i-1}\lessdot\mathcal{G}_{i}$ or $\mathcal{G}_{i-1}\gtrdot\mathcal{G}_{i}$}\right\}.$
We note that the shortest path length is a standard way to measure distance in posets (Monjardet, 1981; Trotter, 1992; Brualdi, Graves and Lawrence, 1995), although its use has primarily been confined to the order theory literature.
Remark 1.
Although a poset always admits a dual poset by replacing $\preceq$’ with $\succeq$’, it is important to adopt and maintain a convention, since some properties are asymmetric. As will become clear in Section˜4, there is, for example, an asymmetry between the empty graph and the full graph, and likewise between lower and upper semimodularity (see Table˜2). This asymmetry is related to the fact that, starting from an arbitrary graph, we can more commonly obtain a graph in the same class by removing edges than adding edges.
3.3 Poset structures that facilitate computing model-oriented distances
Since the model-oriented distance corresponds to the shortest path in the poset’s Hasse diagram, it can, in principle, be computed using search algorithms like A*. However, a naive search is often computationally prohibitive, as it requires repeatedly generating all neighbors for numerous candidate paths. Hence, we investigate structural properties of the poset to save computation — by deriving a closed-form expression for distance, characterizing the neighbors, or pruning the search space. In this subsection, we analyze several such structural properties, laying the groundwork for our specific analysis of probabilistic UGs and CPDAGs, as well as causal DAGs and MPDAGs, in Section˜4.
@@ 3.3.1 Graded poset
A connected poset $\mathcal{L}$ is graded if it admits a function $\operatorname{\mathrm{rank}}:\mathfrak{G}\rightarrow\mathbb{N}$ such that (i) $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ implies $\operatorname{\mathrm{rank}}(\mathcal{G}_{2})=\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1$, and (ii) $\operatorname{\mathrm{rank}}(\mathcal{G})=0$ holds for some minimal element $\mathcal{G}$. It can be shown that if it exists, such a rank function is unique.
@@
Proposition 1 (Distance between comparable graphs in a graded poset).
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is a graded poset with rank function $\operatorname{\mathrm{rank}}(\cdot)$. For any $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ with $\mathcal{G}_{s}\succeq\mathcal{G}_{t}$, every downward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ has length $\operatorname{\mathrm{rank}}(\mathcal{G}_{s})-\operatorname{\mathrm{rank}}(\mathcal{G}_{t})$, which also equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
This is a well-known result that follows from the property that in a graded poset all maximal chains between two comparable elements have equal length. A subset of $\mathfrak{G}$ is a chain (i.e., totally ordered set) if every pair of elements are comparable. A chain is maximal if it is not contained by a larger chain. For completeness, we provide a short proof in Section˜C.1. Gradedness and the rank can greatly facilitate the enumeration of neighborhood, which is the basic routine to any search algorithm. Under gradedness, Eq.˜5 becomes
$\text{Neighbors}(\mathcal{G})=\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\preceq\mathcal{G},\,\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})=\operatorname{\mathrm{rank}}(\mathcal{G})-1\}\\ \cup\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\succeq\mathcal{G},\,\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})=\operatorname{\mathrm{rank}}(\mathcal{G})+1\},\quad\mathcal{G}\in\mathfrak{G}.$
@@ 3.3.2 Down-up and up-down distance, semi-lattice, semimodularity
To reduce the number of paths we need to explore, it is natural to consider the shortest path among a restricted set of paths between two graphs in the Hasse diagram. Specifically, for $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we consider the set of down-up and up-down paths between them, denoted by $\mathrm{P}_{\mathcal{L},\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ and $\mathrm{P}_{\mathcal{L},\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ respectively. A path is down-up if it takes the form of
@@
where $i^{\ast}\in\{0,\dots,d\}$ is the inflection point; downward and upward paths are special cases of down-up paths. Conversely, an up-down path takes the form of
@@
Proposition 2 (Existence of down-up and up-down paths).
The set of down-up paths $\mathrm{P}_{\mathcal{L},\downarrow\uparrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a least element (Condition˜3). Similarly, the set of up-down paths $\mathrm{P}_{\mathcal{L},\uparrow\downarrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a greatest element (Condition˜3).
@@
The posets of all graph types (UGs, DAGs, CPDAGs, and MPDAGs) we consider in this paper have the empty graph as the least element. Further, except for DAGs, all of them also have a greatest element. See Table˜2 for a summary. The down-up and up-down paths lead to the definition of down-up and up-down distances.
@@
Under Condition˜3, the down-up distance is defined as $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L},\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. Under Condition˜3, the up-down distance is defined as $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L},\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
@@
Proposition 3 (Down-up and up-down distances for graded posets).
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\operatorname{\mathrm{rank}}(\cdot)$. Then, under Condition˜3, we have
@@
Any down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ moves downward from $\mathcal{G}_{s}$, and then upward from an inflection point $\mathcal{G}$ to $\mathcal{G}_{t}$. Since $\mathcal{G}\preceq\mathcal{G}_{s}$ and $\mathcal{G}\preceq\mathcal{G}_{t}$, Proposition 1 implies that the length of the path is $\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G})$. Therefore, the shortest down-up path corresponds to an inflection point $\mathcal{G}$ that maximizes $\operatorname{\mathrm{rank}}(\mathcal{G})$ over all $\mathcal{G}\preceq\mathcal{G}_{s},\mathcal{G}\preceq\mathcal{G}_{t}$. The proof for the up-down distance follows similarly. ∎
For a further simplification, we require the notions of meet and join semi-lattice structures: graph $\mathcal{G}_{3}\in\mathfrak{G}$ is the least upper bound (or join) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, denoted as $\mathcal{G}_{1}\vee\mathcal{G}_{2}$, if $\mathcal{G}_{3}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$ and $\mathcal{G}_{3}\preceq\mathcal{G}^{\prime}$ for every $\mathcal{G}^{\prime}$ satisfying $\mathcal{G}^{\prime}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$. The greatest lower bound (or meet) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is defined similarly and is denoted as $\mathcal{G}_{1}\wedge\mathcal{G}_{2}$. Poset $\mathcal{L}$ is a join (meet) semi-lattice if $\mathcal{G}\vee\mathcal{G}^{\prime}$ ($\mathcal{G}\wedge\mathcal{G}^{\prime}$) exists for every $\mathcal{G},\mathcal{G}^{\prime}\in\mathfrak{G}$.
Corollary 1 (Simplification with semi-lattice structure).
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\operatorname{\mathrm{rank}}(\cdot)$. Under Condition˜3, if $\mathcal{L}$ is a meet semi-lattice, we have
@@
We prove Theorem˜2 in Appendix˜D. Fig.˜4 illustrates the basic idea: when $\mathcal{L}$ is semimodular, any shortest path can be transformed into a down-up or up-down path with the same length. We remark that, to our best knowledge, Theorem˜2 weakens the conditions for existing results in the literature and is well-suited for studying model-oriented graph distances. For example, Monjardet (1981, Theorem 8) additionally requires the poset to be graded with a greatest element, while Foldes and Radeleczki (2021, Proposition 2.5) instead requires the poset to be a semi-lattice. We will use Theorem˜2 to study the model-oriented distance for polytree CPDAGs and MPDAGs in Sections˜4.3.2 and 4.4.1.